How old is the bone if if it is presently contains 0.3125 g od C-14, but it was estimated to have originally contained 80.000 g of C-14. the half life of Carbon-14 is 5730 years.


n = ?

T = ? years


1
Expert's answer
2021-04-20T16:49:19-0400

"n=n_0(\frac{1}{2})^{\frac{T}{T_{1/2}}}"


"0.3125g=80.000g\times(\frac{1}{2})^{\frac{T}{5730years}}"


Solving for T,

T = 45840 years


Answer: 45840 years